OpenStudy (anonymous):

-8 + log base 9 (m+1) = -8

OpenStudy (anonymous):

the answer is -1, because you do -8 9(m+1)=-8 then you do nine times m plus nine times one is +-8 so your equation should look like this -8+9m+9=-8, then you do nine minus the eight on the other side of equal sign, and then you get a -17, and then your equation should look like this -8+9m=-17 then you add nine to -17, then your equation is like this 9m=-9, then you divide both sides by -9 , then you get a negative one that is your answer text me back if this is the right answer.